RESONANCE RAMAN SPECTROSCOPY OF DIMETHYL DISULFIDE

\underline{MICHAEL R. WEDLOCK}, , Department of Chemistry, Gettysburg College, Gettysburg, PA 17325.

A resonance Raman spectrum can provide useful information on the initial dissociation dynamics of a molecule as it leaves the Franck-Condon region. We present the resonance Raman spectrum of dimethyl disulfide obtained after excitation at 266 nm. While the literature suggests a prompt dissociation of the S-S bond, the resonance Raman spectrum shows activity in the S-C coordinate.